OpenStudy (anonymous):

What is the distance between the points (-2,1) and (3,7)?

OpenStudy (anonymous):

A\[\sqrt{11}\] B\[\sqrt{37}\] C\[\sqrt{61}\]D\[\sqrt{89}\]

OpenStudy (anonymous):